Find the equation of line passing through (2,2) and having intercepts whose sum is -1

Grade:12

1 Answers

Vikas TU
7 years ago
let the eqn. of the line is:
x/a + y/b = 1
now pass it through (2,2) coordinates.
2/a + 2/b = 1......(1)
and
given
a+b =  1.............(2)
solve these two eqns. and get the valiues of a and b.